OpenStudy (anonymous):

Simplify : v28a^3b^2

OpenStudy (anonymous):

\[\sqrt{28a ^{3}b ^{2}}\]

OpenStudy (anonymous):

I think its 14vab

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{28a^3b^2}\] \[\Large \sqrt{4*7*a^2*a*b^2}\] \[\Large \sqrt{4}*\sqrt{7}*\sqrt{a^2}*\sqrt{a}*\sqrt{b^2}\] \[\Large 2*\sqrt{7}*a*\sqrt{a}*b\] \[\Large 2ab\sqrt{7a}\]